# radarODE: An ODE-Embedded Deep Learning Model for Contactless ECG Reconstruction from Millimeter-Wave Radar

### v1

@@ Abstract
Radar-based contactless cardiac monitoring has become a popular research direction recently, but the fine-grained electrocardiogram (ECG) signal is still hard to reconstruct from millimeter-wave radar signal. The key obstacle is to decouple the cardiac activities in the electrical domain (i.e., ECG) from that in the mechanical domain (i.e., heartbeat), and most existing research only uses pure data-driven methods to map such domain transformation as a black box. Therefore, this work first proposes a signal model for domain transformation, and then a novel deep learning framework called radarODE is designed to fuse the temporal and morphological features extracted from radar signals and generate ECG. In addition, ordinary differential equations are embedded in radarODE as a decoder to provide morphological prior, helping the convergence of the model training and improving the robustness under body movements. After being validated on the dataset, the proposed radarODE achieves better performance compared with the benchmark in terms of missed detection rate, root mean square error, Pearson correlation coefficient with the improvement of $9\%$, $16\%$ and $19\%$, respectively. The validation results imply that radarODE is capable of recovering ECG signals from radar signals with high fidelity and can be potentially implemented in real-life scenarios.
@@ I Introduction
Radar-based human sensing is a rapidly evolving field that leverages radio-frequency signals to detect or recognize human activities in many future scenarios (e.g., smart home and in-cabin monitoring [1, 2, 3]). Compared with other contactless sensors such as cameras and acoustic sensors with privacy issues, radar could sense the ambient environment in a non-invasive manner and achieve good robustness under light conditions or temperature variations [4, 5]. In exchange, radar signals cannot be directly interpreted by humans like other mediums with explicit meanings (e.g., images, sound), increasing the complexity of designing efficient frameworks for specific sensing tasks (e.g., pose estimation [6], object detection [7] and vital sign monitoring [8]).
@@
Within all scenarios suitable for radar-based human sensing, contactless vital sign monitoring is a crucial task in providing healthcare information (e.g., respiration, heart rate and electrocardiogram (ECG)). The earliest attempt for radar-based respiration monitoring can be traced back to 1975 by measuring the chest wall displacement induced by respiration [9]. The chest wall displacement will modulate the phase component of the emitted radar signal, and the latent respiration information can be demodulated from the phase variation [10]. Similarly, cardiac activities are small-scale displacements that also cause chest wall displacements, but such small displacements are normally ruined by respiration with orders more amplitude. The follow-up researchers are dedicated to extracting cardiac information in the presence of the respiration disturbance and also other common noises, such as random body movement (RBM) [11, 12], multi-path or multi-person interference [13, 14] and radar self-movement [15, 16].
@@
In the context of cardiac monitoring, most early studies focused on the recovery of coarse cardiac information, such as heart rate (HR), heart sound and heart rate variability (HRV), from the perspectives of radar front-end design or advanced algorithms design [8]. For example, some advanced types of radar (e.g., frequency modulated continuous wave (FMCW) radar) are designed to enable high range-resolution or multi-person monitoring [17], and some baseband signal-processing algorithms are embedded on the radar platform to realize in-phase/quadrature modulation or accurate phase unwrapping [18]. In addition, various advanced algorithms are applied by leveraging different intrinsic characteristics of cardiac activities to robustly reconstruct cardiac features. For example, cardiac activities normally reveal strong periodicity in the time domain and have dominant peaks on the spectrum, inspiring periodicity-based methods (e.g., template matching [19, 20], hidden Markov model [21, 22]) and spectrum-based methods (e.g., Fourier transform [23, 24], wavelet transform [25, 7]) as two major categories in cardiac feature extraction algorithms.
@@
In recent years, the emergence of commercial radar platforms with high operating frequency (millimeter-wave (mm-wave) radar) encourages researchers to extract fine-grained cardiac features (e.g., ECG and seismocardiography (SCG)) from the radar signal [8]. SCG signal is measured by the accelerometer mounted on the human chest to describe the mechanical vibrations produced by heartbeats, describing the fine-grained cardiac mechanical activities such as aortic/mitral valve opening/closing and isovolumetric contraction [26]. Although these vibrations are subtle, it is still reasonable to directly map the displacements detected by radar to each fine-grained cardiac mechanical activity using high-resolution radar as proved in [17]. However, radar-based SCG recovery is not widely investigated compared with ECG, because ECG provides more comprehensive information (e.g., atrial/ventricular depolarization [27]) for clinical diagnosis.
@@
To reconstruct ECG from radar signal, the most straightforward approach is to directly sense the scattered electromagnetic field variation through frequency shift of mm-wave response, and the ECG signal can then be decoupled from the scattered electromagnetic field based on the dynamic model in the form of partial differential equations deduced from cardiac electrophysiology (i.e., ionic concentration in cardiac cells) [28, 29]. However, the solutions of the entire model are extremely hard to obtain either numerically or analytically, and the constructed models will be changed with respect to different environments and noises due to the Green’s function [30], causing difficulty in adapting the model in various real-life scenarios.
@@
The second approach, which is also the most adopted approach, only uses radar to sense the cardiac mechanical activities through the phase-modulated mm-wave response as been investigated in coarse cardiac monitoring, but then the researchers must deal with domain decoupling to transform the measured signal from the mechanical domain to the electrical domain to generate the final ECG measurement. Intuitively, it is reasonable to say that mechanical conduction and electrical conduction are highly correlated in describing cardiac activities because the electrical changes in cells trigger the heart muscle contraction, whereas such a relationship is called excitation-contraction coupling in electrophysiology and is extremely hard to interpret or model by the people without biological knowledge [31, 32].
@@
In literature, the existing studies all leverage deep learning methods to extract the latent information from enormous training data and try to learn the domain transformation relying on the extraordinary non-linear mapping ability of the deep neural network [33, 34, 35, 36]. Although these studies could successfully reconstruct the ECG signal from radar, three drawbacks still need to be improved:
@@
  • There is no existing signal model with an easily understood form that can be integrated into domain transformation for radar-based ECG reconstruction.
@@
  • The pure data-driven method could learn the domain transformation as a black box, but the researchers can hardly intervene in the learning process to enhance the characteristic peaks for ECG or provide prior knowledge to help with network training.
@@
  • The well-trained deep learning model is not robust to abrupt noises such as body movement [35, 37], because these noises normally have orders of magnitude larger than cardiac-related vibrations, drowning the subtle features and ruining the forward propagation of the deep neural network [38].
@@
Inspired by the above discussions, this paper aims to design a framework for radar-based ECG reconstruction with ordinary differential equations (ODEs) embedded to provide prior knowledge on domain transformation. The contributions of this research can be concluded as:
@@
  • This study innovatively models the radar signal in terms of fine-grained cardiac features, enabling further transformation between cardiac mechanical and electrical activities, instead of using pure data-driven approaches as in the literature.
@@
  • Based on the proposed signal model, an ODE-embedded module called single cycle ECG generator (SCEG) is designed to parameterize the radar signal into sparse representations with morphological meanings, enhancing the characteristic peaks of the recovered ECG signal and also accelerating the convergence of the model training process.
  • By fusing the extracted morphological and temporal features, the proposed radarODE framework is proven to be robust in the presence of body movement noise and could realize accurate reconstruction of the ECG signal from the millimeter wave radar signal.
@@
The rest of the paper is organized as follows. Section II introduces the background knowledge required for radar-based ECG reconstruction. Section III explains the proposed model for radar signal and the structure of radarODE framework. Section IV first introduces the public dataset used for validation in this research and then presents the results obtained with corresponding comparisons and evaluations. Finally, Section V concludes this paper.
@@ II Background
To understand the model and framework proposed later in this paper, this section will first provide the necessary background about the signal model for radar-based coarse cardiac monitoring and then briefly introduce the relationship between the cardiac electrical and mechanical activities.
@@ II-A Radar-Based Coarse Cardiac Monitoring
where $\lambda$ is the wavelength that equals to $\frac{c}{f}$. According to [39, 40], it is safe to eliminate the changes in amplitude and phase noise term because the chest displacement is much less than the fixed distance (i.e., $x(t)\ll d_{0}$). Therefore, the approximate received signal is
@@
where $\theta_{d}$, $\theta_{0}$ and $\Delta\theta(t)$ are phase shifts affected by different factors such as $d_{0}$, signal mixer and antenna, and can be set as constant in this paper [40]. Then, the phase signal unwrapped from the baseband signal is obtained as
@@
The follow-up researchers have proposed various techniques to improve the accuracy of the unwrapped phase signal variation in (9). For example, the in-phase/quadrature modulation is proposed to solve the null point issue [39]; the differentiate and cross-multiply algorithm is designed to avoid discontinuity in the unwrapped phase signal [41]. In addition, chest displacement $x(t)$ is a mixture of cardiac activities, respiration and noises (e.g., RBM [11, 12], multi-path or multi-person interference [13, 14, 42]). Therefore, enormous advanced algorithms are proposed to decompose the cardiac information from $x(t)$, as have been reviewed in [8].
II-B Radar-Based Fine-Grained Cardiac Monitoring
The coarse cardiac monitoring only aims to detect a single heartbeat within one cardiac cycle, and different cardiac features (e.g., HR, BBI and HRV) can then be calculated through the time indexes of a series of heartbeats. In contrast, fine-grained cardiac monitoring requires recovering as many cardiac activities as possible within one cardiac cycle. For example, Figure 1 shows the typical ECG and SCG signal waveform, representing the cardiac electrical and mechanical activities, respectively. ECG waveform comprises P-wave, QRS-complex and P-wave, corresponding to atrial depolarization, ventricular depolarization and ventricular repolarization, respectively. These electrical changes trigger the heart muscle contraction, causing the aortic valve opening/closure (AO/AC) and mitral valve opening/closure (MO/MC) [31, 32], as shown in the SCG plot in Figure 1. In addition, The relationship between cardiac electrical and mechanical activities is called excitation–contraction coupling in electrophysiology and can be modeled mathematically according to the intracellular calcium concentration, because calcium transient is the link between the action potential and muscle contraction [31, 32].
Figure 1: Single cycle of ECG and SCG signals with corresponding cardiac activities and radar signal in the public dataset [35]
In the context of ECG reconstruction, two types of approaches in literature have been generalized in the Introduction section. The first approach is not considered in this paper because it involves complex modeling of electrophysiology from biology and chemistry [28]. This research only focuses on the second approach by reconstructing the ECG from the mechanical activities measured by radar. Although the underlying principle of the second approach is exactly the same as derived in (9), the radar signal should have a higher signal-to-noise ratio (SNR) and resolution than that used for coarse monitoring. For example, the mechanical activities can be simplified as the first vibration related to AO and the second vibration related to AC [17], corresponding to the QRS-complex and T-wave in the ECG signal. Therefore, the radar signal with high SNR and resolution should be able to reveal these two vibrations $v_{1}$ and $v_{2}$ as shown in the radar signal in Figure 1. In this case, only one public dataset [35] collected by $77$ GHz FMCW radar satisfies the requirement and complies with our proposed model in the next section.
@@ III-A Overview
Figure 2: Overview of the radarODE framework with three modules: (a) Single Cycle ECG Length Estimation; (b) Single Cycle ECG Generator; (c) Long-term ECG Reconstruction.
In order to address the drawbacks in radar-based ECG recovery as mentioned in the Introduction, this section will first introduce the proposed signal model that considers the nature of the fine-grained cardiac mechanical activities, ensuring a reasonable transformation towards cardiac electrical activities (ECG) instead of using pure data-driven methods as a black box. Afterwards, the proposed radarODE framework will be elaborated to fulfill the transformation from radar to ECG signal with $50$ synchronous radar signals [35] as inputs, representing the measurements from $50$ spatial points within the chest region as shown in Figure 2. The radarODE is composed of three modules to output the final ECG reconstructions as:
  • The first module estimates the peak-to-peak interval (PPI) for consecutive heartbeats from the energy plot of the input radar signals based on the proposed peak detection algorithm, as shown in Figure 2(a).
  • The second module first transforms the time-domain radar signal into spectrograms using synchrosqueezed wavelet transform (SST) to reveal the latent information in the frequency domain. Then, the morphological features (i.e., amplitudes and positions of PQRST peaks) for the single cycle ECG signal can be extracted from SST plots using the proposed SCEG, as shown in Figure 2(b).
  • The third module is used to fuse the extracted morphological features with the temporal features hidden in the original radar signal to generate the final ECG recovery, as shown in Figure 2(c).
@@ III-B1 Fine-Grained Model for Radar Signal
Furthermore, the pre-processed radar signal $\tilde{x}(t)$ has two prominent vibrations $v_{1}$ and $v_{2}$ as shown in Figure 1, corresponding to the fine-grained cardiac mechanical activities shown in SCG. According to the previous work on SCG modeling, the heart muscle contraction has a pulsatile nature [23], and the bones/tissues in chest area introduce the extra damping into the pulse [43]. Inspired by the natural characteristics, the radar signal with two prominent vibrations measured in a single cardiac cycle is innovatively modeled as the Gaussian pulses with certain central frequencies as
@@
The aim of proposing the model in (12) is not to perform the curve fitting since the goal of this paper is ECG reconstruction but not SCG. In fact, the time indices of the vibrations ($T_{1}$ and $T_{2}$) are critical and need to be identified to help reconstruct QRS-complex and T-wave in ECG signals. In addition, to verify the fidelity of the proposed model in (12), a piece of high SNR radar signal $\tilde{x}(t)$ with the corresponding synthesis is shown in Figure 3. Empirically, the central frequencies of $v_{1}$ and $v_{2}$ are laid in the range of $[1,15]$ Hz and $[20,25]$ Hz respectively, and all the unknown parameters can be obtained by solving the linear regression problem with respect to the band-passed radar signal. For example, Figure 3(a) shows the radar signal $\tilde{x}(t)$ filtered by the bandpass filter (BPF) with range $[1,15]$ Hz, and the corresponding synthesis $v_{1}$ is obtained as $a_{1}=0.22$, $b_{1}=0.05$, $f_{1}=10$ and $T_{1}=-0.69$. Similarly, the synthesis of the radar signal after BPF ($[20,25]$ Hz) is shown in Figure 3(b) with $a_{2}=0.12$, $b_{2}=0.03$, $f_{2}=23$ and $T_{2}=-0.33$. The overall result $v_{1}+v_{2}$ shows that the synthesis based on the proposed model fits the original radar signal well, and the time indices of two vibrations can be found accurately.
[…]
Remark 1. Although the signal model proposed in this section is only used for the alignment between the cardiac mechanical and electrical features, this model actually describes the fundamental cardiac activities within a single cardiac cycle in a more realistic manner than only considering the single-tone heartbeat. Therefore, the proposed model is potentially universal to many other cardiac-related tasks as a template, such as heart rate monitoring, cardiac events segmentation and synthesis radar measurement generation/augmentation.

### v2

@@ Abstract
Radar-based cardiac monitoring has become a popular research direction recently, but the fine-grained electrocardiogram (ECG) signal is still hard to reconstruct from millimeter-wave radar signal. The key obstacle is to decouple cardiac activities in the electrical domain (i.e., ECG) from that in the mechanical domain (i.e., heartbeat), and most existing research only uses purely data-driven methods to map such domain transformation as a black box. Therefore, this work first proposes a signal model that considers the fine-grained cardiac feature sensed by radar, and a novel deep learning framework called radarODE is designed to extract both temporal and morphological features for generating ECG. In addition, ordinary differential equations are embedded in radarODE as a decoder to provide morphological prior, helping the convergence of the model training and improving the robustness under body movements. After being validated on the dataset, the proposed radarODE achieves better performance compared with the benchmark in terms of missed detection rate, root mean square error, Pearson correlation coefficient with improvements of $9\%$, $16\%$ and $19\%$, respectively. The validation results imply that radarODE is capable of recovering ECG signals from radar signals with high fidelity and can potentially be implemented in real-life scenarios.
@@ I Introduction
Radar-based human sensing is a rapidly evolving field that leverages radio-frequency signals to detect or recognize human activities in many future scenarios (e.g., smart home, in-cabin monitoring and biometric recognition [1, 2, 3]). Compared with other contactless sensors such as cameras and acoustic sensors with privacy issues, radar could sense the ambient environment in a non-invasive manner and achieve good robustness under light conditions or temperature variations [4, 5]. In exchange, radar signals cannot be directly interpreted by humans like other mediums with explicit meanings (e.g., images, sound), increasing the complexity of designing efficient frameworks for specific sensing tasks (e.g., pose estimation [6], object detection [7] and vital sign monitoring [8]).
@@
Within all scenarios suitable for radar-based human sensing, contactless vital sign monitoring is a crucial task in providing healthcare information (e.g., respiration, heart rate and electrocardiogram (ECG)). The first attempt at radar-based respiration monitoring can be traced back to 1975 by measuring the displacement of the chest wall induced by respiration [9]. The chest wall displacement will modulate the phase component of the emitted radar signal, and the latent respiratory information can be demodulated from the phase variation [10]. Similarly, cardiac activities are small-scale displacements that also cause chest wall displacements, but such small displacements are normally ruined by respiration with orders more amplitude. The follow-up researchers are dedicated to extracting cardiac information in the presence of respiration disturbance and also other common noises, such as random body movement (RBM) [11, 12], multi-path or multi-person interference [13, 14] and radar self-movement [15, 16].
@@
In the context of cardiac monitoring, most early studies focused on the recovery of coarse cardiac information, such as heart rate (HR), heart sound and heart rate variability (HRV), from the perspectives of radar front-end design or advanced algorithms design [8]. For example, some advanced types of radar (e.g., frequency modulated continuous wave (FMCW) radar) are designed to enable high range-resolution or multi-person monitoring [17], and some baseband signal-processing algorithms are embedded on the radar platform to realize in-phase/quadrature modulation or accurate phase unwrapping [18]. In addition, various advanced algorithms are applied by leveraging different intrinsic characteristics of cardiac activities to robustly reconstruct cardiac features. For example, cardiac activities normally reveal strong periodicity in the time domain and have dominant peaks on the spectrum, inspiring periodicity-based methods (e.g., template matching [19], hidden Markov model [20]) and spectrum-based methods (e.g., Fourier transform [21], wavelet transform [7]) as two major categories in cardiac feature extraction algorithms.
@@
In recent years, the emergence of commercial radar platforms with high operating frequency (millimeter-wave (mm-wave) radar) encourages researchers to extract fine-grained cardiac features (e.g., ECG and seismocardiography (SCG)) from the radar signal [8]. SCG signal is measured by the accelerometer mounted on the human chest to measure the mechanical vibrations produced by heartbeats, describing the fine-grained cardiac mechanical activities such as aortic/mitral valve opening/closing and isovolumetric contraction [22]. Although these vibrations are subtle, it is still reasonable to directly map the displacements detected by radar to each fine-grained cardiac mechanical activity using high-resolution radar as proved in [17]. However, radar-based SCG recovery is not widely investigated compared with ECG, because ECG provides more comprehensive information (e.g., atrial/ventricular depolarization [23]) for clinical diagnosis.
@@
To reconstruct ECG from radar signal, the most straightforward approach is to directly sense the variation in the scattered electromagnetic field through frequency shift of mm-wave response, and the ECG signal can then be decoupled from the scattered electromagnetic field based on the dynamic model in the form of partial differential equations deduced from cardiac electrophysiology (i.e., ionic concentration in cardiac cells) [24]. However, the solutions of the entire model are extremely hard to obtain either numerically or analytically, and the constructed models will be changed with respect to different environments and noises due to the Green’s function [25], causing difficulty in adapting the model in various real-life scenarios.
@@
The second approach, which is also the most adopted approach, only uses radar to sense the chest region displacement through the reflected radar signal as in coarse cardiac monitoring, but then the researchers must deal with domain decoupling to transform the measured signal from the mechanical domain to the electrical domain to generate ECG measurement. Intuitively, it is reasonable that mechanical conduction and electrical conduction are highly correlated in describing cardiac activities because the electrical changes in cells trigger heart muscle contraction, whereas such a relationship is called excitation-contraction coupling in electrophysiology and is extremely hard to interpret or model by researchers without biological knowledge [26, 27].
@@
In the literature, the existing studies all leverage deep learning methods to extract latent information from enormous radar/ECG pairs and try to learn domain transformation relying on the extraordinary non-linear mapping ability of the deep neural network [28, 29, 30, 31, 32]. Although these studies could successfully reconstruct the ECG signal from radar, three drawbacks still need to be improved:
@@
  • There is no existing signal model with a compact form to describe the domain transformation for radar-based ECG reconstruction.
@@
  • The purely data-driven method could learn the domain transformation as a black box, but researchers can hardly intervene in the learning process to enhance the characteristic peaks of ECG or explain the intrinsic correspondence in domain transformation.
@@
  • The well-trained deep learning model is not robust to abrupt noises such as body movement [28, 33], because these noises normally have orders of magnitude higher than cardiac-related vibrations, drowning out subtle features and ruining forward propagation of the deep neural network [34].
@@
Inspired by the above discussions, this paper aims to design a framework for radar-based ECG reconstruction with ordinary differential equations (ODEs) embedded to provide prior knowledge on domain transformation and resist abrupt noises. The contributions of this research can be concluded as:
  • This study proposes an ODE-embedded framework called radarODE to produce robust long-term ECG recovery against abrupt noises with the aid of morphological ECG features as the reference.
@@
  • A signal model is designed to describe the fine-grained cardiac feature sensed by radar, enabling further domain transformation between cardiac mechanical and electrical activities, instead of using purely data-driven approaches without any explanation as in the literature.
@@
  • Based on the proposed signal model, an ODE-embedded module called single-cycle ECG generator (SCEG) is designed to realize the domain transformation by parameterizing the radar signal into sparse representations and hence generate the morphological ECG features as references to resist noise.
@@
The rest of the paper is organized as follows. Section II introduces the background knowledge required for radar-based ECG reconstruction. Section III explains the proposed model for radar signal and the structure of radarODE framework. Section V and IV introduce the public dataset used for validation in this research and then present the results obtained with corresponding comparisons and evaluations. Finally, Section VI concludes this paper.
@@ II Background
To understand the model and framework proposed later in this paper, this section will first provide the necessary background about the signal model for radar-based coarse cardiac monitoring and then briefly introduce the relationship between cardiac electrical and mechanical activities.
@@ II-A Radar-Based Coarse Cardiac Monitoring
where $\lambda$ is the wavelength that equals to $\frac{c}{f}$. According to [35, 36], it is safe to eliminate changes in amplitude and phase noise term because the chest displacement is much less than the fixed distance (i.e., $x(t)\ll d_{0}$). Therefore, the approximate received signal is
@@
where $\theta_{d}$, $\theta_{0}$ and $\Delta\theta(t)$ are phase shifts affected by different factors such as $d_{0}$, signal mixer and antenna, and can be set as constant [36]. Then, the phase signal unwrapped from the baseband signal is obtained as
@@
The follow-up researchers have proposed various techniques to improve the accuracy of the unwrapped phase signal variation in (9). For example, the in-phase/quadrature modulation is proposed to solve the null point issue [35]; the differentiate and cross-multiply algorithm is designed to avoid discontinuity in the unwrapped phase signal [8]. In addition, chest displacement $x(t)$ is a mixture of cardiac activities, respiration and noises (e.g., RBM [11, 12], multi-path or multi-person interference [13, 14]). Therefore, enormous advanced algorithms are proposed to decompose cardiac information from $x(t)$, as have been reviewed in [8].
II-B Radar-Based ECG Monitoring as a Domain Transformation Problem
Coarse cardiac monitoring only aims to detect a single heartbeat within one cardiac cycle, while fine-grained cardiac monitoring requires recovering subtle cardiac activities within one cardiac cycle. For example, Figure 1 shows the typical radar and SCG signal waveform within a single cardiac cycle that describes the cardiac mechanical activities, such as aortic valve opening/closure (AO/AC) and mitral valve opening/closure (MO/MC) [26, 27]. These mechanical activities are muscle contractions stimulated by cardiac electrical events, such as P-wave, QRS-complex and P-wave in the ECG signal as shown in Figure 1. Therefore, radar-based ECG recovery is actually a domain transformation problem that translates cardiac mechanical activities into electrical activities and realizes fine-grained vital sign monitoring in a contactless manner.
In the literature, the domain transformation is only realized by deep-learning-based methods, while a common issue of the deep learning model is not robust against large-scale noise (e.g., RBM) as reported by many previous studies [29, 30, 31, 32]. However, there is no existing work that investigates the noise-robustness of the deep learning model itself, and this study is motivated to develop a noise-robust deep learning model to realize the domain transformation in the presence of body movement noise.
Figure 1: Relationships between cardiac mechanical and electrical activities within the same cardiac cycle.
@@ III-A Overview
Figure 2: Overview of the radarODE framework for domain transformation: (a) Single-cycle ECG length (PPI) estimation; (b) Single-cycle ECG generator; (c) Long-term ECG reconstruction.
In order to realize a robust radar-to-ECG reconstruction, the proposed radarODE framework is designed based on the domain transformation within a single cardiac cycle to generate faithful ECG reference that aids the domain transformation in long-term ECG recovery. The inputs of radarODE are $50$ synchronous radar signals that represent the measurements from $50$ spatial points within the chest region [28], as shown in Figure 2, and the entire domain transformation is realized by three modules:
  • The first module estimates the peak-to-peak interval (PPI) for consecutive heartbeats based on the proposed peak detection algorithm and provides information for single cardiac cycle segmentation, as shown in Figure 2(a).
  • The second module first transforms the time-domain radar signal within a single cardiac cycle into a spectrogram using synchrosqueezed wavelet transform (SST). Then, the temporal and ODE decoder within SCEG will generate the faithful ECG pieces as morphological references while resisting noise disruption, as shown in Figure 2(b).
  • The third module is used to fuse the extracted morphological and temporal features hidden in the original radar signal to generate the final ECG recovery, as shown in Figure 2(c).
@@ III-B1 Fine-Grained Model for Radar Signal
Furthermore, the pre-processed radar signal $\tilde{x}(t)$ has two prominent vibrations $v_{1}$ and $v_{2}$ as shown in Figure 1, corresponding to the fine-grained cardiac mechanical activities shown in SCG. According to the previous work on SCG modeling, the heart muscle contraction has a pulsatile nature, and the bones/tissues in chest area introduce the extra damping into the pulse [21]. Inspired by the natural characteristics, the radar signal with two prominent vibrations measured in a single cardiac cycle is innovatively modeled as the Gaussian pulses with certain central frequencies as
@@
The aim of proposing the model in (12) is not to perform the curve fitting but to provide the explanation for the later radarODE design, because the positions of the prominent vibrations ($T_{1}$, $T_{2}$) are crucial to the precise reconstruction of ECG peaks using deep neural network.